The maximum value of (logx)/x is
  1. 2/e
  2. 1/e
  3. 1
  4. e
Please explain it.

Grade:Select Grade

2 Answers

noogler
8 years ago
1/e
let f(x)=logx/x
to get max of f(x)............ fi(x)=0 u get some value of x
substitute it in f(x) u get max value of f(x)
Yash Chourasiya
3 years ago
Dear Student

Let f = logx/x

Derivative is
df/dx ​= (1−logx)/x2​
maximum value occurs when derivative is zero.
So⇒ (1 − logx)/x2​ = 0
⇒logx = 1
⇒x = e

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya